OpenStudy (anonymous):

If y = 2 when x = –4, find x when y = 20.

OpenStudy (anonymous):

-40

OpenStudy (texaschic101):

2/-4 = 20/x cross multiply 2x = (20)(-4) 2x = -80 x = -40

OpenStudy (arabpride):

set up something like a proportion: so if y = 2(x) = 20 x would = 10 and if you multiply -4 x 10 ur answer is -40

OpenStudy (anonymous):

If y = 36 when x = 9, find x when y = 52

OpenStudy (anonymous):

04 / 2 = -2. 20 x -2 = -40. Thus, x = -40 when y = 20

OpenStudy (anonymous):

*-4

OpenStudy (texaschic101):

36/9 = 52/x cross multiply 36x = (52)(9) 36x = 468 x = 468/36 x = 13

OpenStudy (anonymous):

@texaschic101 these are the possible answers a. y = –4x; –13 c. y = 4x; 10.4 b. y = 3x; 13 d. y = 4x; 13

OpenStudy (texaschic101):

D. y = 4x ; 13